OpenStudy (anonymous):

EVALUATE THE PRINCIPAL SQUARE ROOT OF 36 .???

OpenStudy (saifoo.khan):

\[\Large \sqrt{ 36} \to \sqrt{6 \times 6} = 6\]

OpenStudy (anonymous):

Principal square root is only the positive sq rt